Documenta Mathematica

Band 20, 2015

Mathieu Huruguen Log Homogeneous Compactifications

of Some Classical Groups  1–35

Hubert Kalf, Takashi Okaji, and Osanobu Yamada The Dirac Operator with Mass m_{0} ≥ 0 : Non-Existence of Zero Modes

and of Threshold Eigenvalues  37–64

Giulia Battiston The Variation of the Monodromy Group in Families

of Stratified Bundles in Positive Characteristic  65–87

iv